OpenStudy (anonymous):

Find the number of ways in which three numbers can be selected from the set \(\{1,2,3,\ldots,4n\}\), such that the sum of the three selected numbers is divisible by 4.

OpenStudy (anonymous):

1+2+3=4n is this ?

OpenStudy (anonymous):

type.

OpenStudy (anonymous):

anyway whatever :p

OpenStudy (experimentx):

n numbers divisible by 4, nC3 <--- perhaps this will be for all three numbers divisible by 4 on the other hand, we we choose odd number, say 1, then we have to choose 3 or 7, 5 won't work, I think there are n such numbers, n*(n-1)*n so total wouldbe nC3 + n*(n-1)*n <--- just a guess

OpenStudy (anonymous):

How about we express every number as \(4n,4n+1,4n+2\) and \(4n+3\) ? Case1. 4n, 4n, 4n Case2. 4n, 4n + 1, 4n + 3 Case3. 4n + 2, 4n + 2, 4n Case4. 4n+1,4n+1,4n+2 Would this work?

OpenStudy (experimentx):

hey man, can you choose the same no twice??

OpenStudy (anonymous):

No? it's just the form of number, not the number itself.

OpenStudy (experimentx):

Case4. 4n+1,4n+1,4n+2 <-- not sure about this

OpenStudy (experimentx):

looks like i missed case 3

OpenStudy (anonymous):

Since there are 4n numbers, each form must have n numbers. For Case1. n*(n-1)*(n-2) Case2. n*n*n Case3. n*(n-1)*n Case4. n*(n-1)*n

OpenStudy (experimentx):

is it combination or arrangement??

OpenStudy (anonymous):

experimentx just the form of number. For example \(\underbrace{1}_{4\cdot0 +1} + \underbrace{5}_{4\cdot1 + 1} + \underbrace{2}_{4\cdot 0 +2} = 8\).

OpenStudy (anonymous):

Hmm so total number of ways must be n^3 + n(n-1)(n-2) + 2(n)*(n-1)*n

OpenStudy (experimentx):

seems all right to me.

OpenStudy (anonymous):

But I can't be sure, I don't have the answer. :(

OpenStudy (experimentx):

For Case1. n*(n-1)*(n-2) <--- same pattern so it should be decreasing Case2. n*n*n <--- three different pattern Case3. n*(n-1)*n <---- two same pattern one different pattern Case4. n*(n-1)*n <--- two same pattern one different patter

OpenStudy (experimentx):

if it is arrangement, then it should be that way. if it's just selection, then we have to divide it by 3!

OpenStudy (anonymous):

ohh yeah, it's selection. so divide it by 3!.

OpenStudy (experimentx):

i think so.

OpenStudy (anonymous):

hmm \[C_1\cdot C_1 \cdot C_1 + 2\cdot C_2\cdot C_1 +C_3\]Is this right? nCr -> Cr.

OpenStudy (experimentx):

not really, 1) nC3 2) nC1 x nC1 x nC1 3) nC2 x nC1 4) nC2 x nC1 I think it would be this way, for all.

OpenStudy (experimentx):

Oo .. I come to think now, dividing by 3! was only in the first case. I would have been incorrect.

OpenStudy (anonymous):

hmm thanks :-)

OpenStudy (experimentx):

well ... most of the job you did yourself. still ... you are welcome.